The Happenstance bakery store is well known for its Canadian style butter tarts. A recent analyst...

  1. The Happenstance bakery store is well known for its Canadian style butter tarts. A recent analyst determined that the store sells on average 18 butter tarts per hour. The store is open from 7 AM to 7PM every day. The analyst also tracked inventory hour-by-hour for the past month and determined that the store averaged being out-of-stock 1.5 hours per day.   What is the fill rate for this bakery for butter tarts? (1)
Answer #1

Given : Store is open from 7 am to 7 pm

Total working hours in a day =12 hours

Total out of stock hours in a day =1.5

Average tarts sold per hour =18

Fill rate refers to the fraction of demand that is satisfied from the product in the inventory without loss of sales or out of stock.

Total number of hours product is in stock=Total working hours -total out of stock hours

=12-1.5

=10.5 hours

Fill rate =(Total demand satisfied in 10.5 hours /Total demand in 12 hours) *100

=(10.5 *18/ 12 *18) *100

=(10.5/12)*100

=87.5 %